# CEReBrO: Compact Encoder for Representations of Brain Oscillations Using Efficient Alternating Attention

### v3

@@ 4.5 Comparisons with the State of the Art
| Methods | Model Size | Balanced Accuracy (%) | AUPR | AUROC |
| Supervised Models | ||||
| SPaRCNet [23] | 0.79M | 78.96 0.18 | 0.8414 0.0018 | 0.8676 0.0012 |
| CNN-Transformer [39] | 3.2M | 77.77 0.22 | 0.8433 0.0039 | 0.8461 0.0013 |
| FFCL [28] | 2.4M | 78.48 0.38 | 0.8448 0.0065 | 0.8569 0.0051 |
| ST-Transformer [45] | 3.5M | 79.66 0.23 | 0.8521 0.0026 | 0.8707 0.0019 |
| Self-supervised Models | ||||
| BrainBERT [52] | 43.18M | - | 0.8460 0.0030 | 0.8530 0.0020 |
| EEGFormer Small [7] | - | - | 0.8620 0.0050 | 0.8620 0.0070 |
| EEGFormer Base [7] | - | - | 0.8670 0.0020 | 0.8650 0.0010 |
| EEGFormer Large [7] | - | - | 0.8720 0.0010 | 0.8760 0.0030 |
| LaBraM [21] | 5.8M | 81.40 0.19 | 0.8965 0.0016 | 0.9022 0.0009 |
| LaBraM [21] | 46M | 82.26 0.15 | 0.9130 0.0005 | 0.9127 0.0005 |
| LaBraM [21] | 369M | 82.58 0.11 | 0.9204 0.0011 | 0.9162 0.0016 |
| CEReBrO | 3.58 M | 79.40 0.19 | 0.8763 0.0031 | 0.8749 0.0033 |
| CEReBrO | 39.95 M | 81.29 0.15 | 0.8994 0.0002 | 0.8867 0.0006 |
@@ 6 Acknowledgements
This work was supported by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies.

### v4

@@ 4.5 Comparisons with the State of the Art
| Methods | Model Size |
AUPR | AUROC | ||
| Supervised Models | ||||||
| SPaRCNet [23] | 0.79M | 78.96 0.18 | 0.8414 0.0018 | 0.8676 0.0012 | ||
| CNN-Transformer [39] | 3.2M | 77.77 0.22 | 0.8433 0.0039 | 0.8461 0.0013 | ||
| FFCL [28] | 2.4M | 78.48 0.38 | 0.8448 0.0065 | 0.8569 0.0051 | ||
| ST-Transformer [45] | 3.5M | 79.66 0.23 | 0.8521 0.0026 | 0.8707 0.0019 | ||
| Self-supervised Models | ||||||
| BrainBERT [52] | 43.18M | - | 0.8460 0.0030 | 0.8530 0.0020 | ||
| EEGFormer Small [7] | - | - | 0.8620 0.0050 | 0.8620 0.0070 | ||
| EEGFormer Base [7] | - | - | 0.8670 0.0020 | 0.8650 0.0010 | ||
| EEGFormer Large [7] | - | - | 0.8720 0.0010 | 0.8760 0.0030 | ||
| LaBraM [21] | 5.8M | 81.40 0.19 | 0.8965 0.0016 | 0.9022 0.0009 | ||
| LaBraM [21] | 46M | 82.26 0.15 | 0.9130 0.0005 | 0.9127 0.0005 | ||
| LaBraM [21] | 369M | 82.58 0.11 | 0.9204 0.0011 | 0.9162 0.0016 | ||
| CEReBrO | 3.58 M | 79.40 0.19 | 0.8763 0.0031 | 0.8749 0.0033 | ||
| CEReBrO | 39.95 M | 81.29 0.15 | 0.8994 0.0002 | 0.8867 0.0006 | ||
@@ 6 Acknowledgements
This work was supported by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies. This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID lp12 on Alps.